# Bayesian reconstruction of primordial perturbations from induced gravitational waves

### v2

@@ 1 Introduction
The power spectrum of curvature perturbations, $\mathcal{P}_{\zeta}(k)$, plays a crucial role in understanding the early Universe and the formation of cosmic structure. While the cosmic microwave background (CMB) and large-scale structure of the universe provide constraints on its features at large cosmological scales (see e.g. [1]), smaller scales remain relatively unconstrained. One promising avenue to probe this regime is through scalar-induced gravitational waves (SIGWs), which arise as a second-order effect of primordial density perturbations: see e.g. [2, 3, 4, 5, 6, 7, 8, 9, 10, 11] for original and key papers, and [12] for a comprehensive, detailed review. Such induced gravitational waves carry valuable information about the small-scale features of $\mathcal{P}_{\zeta}(k)$ and may be detected by future gravitational wave observatories, such as the Laser Interferometer Space Antenna (LISA) [13] and the Einstein Telescope (ET) [14]. Enhanced scalar fluctuations, leading to observable GW signals, can be produced on models leading to primordial black holes, see e.g. [15] for a review. In addition to encoding information about the primordial power spectrum, the gravitational wave (GW) spectrum is also sensitive to the expansion history of the Universe (see e.g. [16, 17, 18]). This is good news since deviations from a purely radiation-dominated phase at very early times can leave distinct imprints on the shape and amplitude of the GW spectrum, which can be inferred from data. Such departures might arise from exotic physics or alternative cosmological scenarios, and they can be parametrized in terms of an effective equation of state parameter [12]. A detection of SIGWs, therefore, offers a unique opportunity not only to reconstruct $\mathcal{P}_{\zeta}(k)$, but also to infer the background expansion dynamics.
@@ 2 Induced GWs and the Universe’s expansion history
In this section we collect and outline the key equations and conventions relevant for the computation of scalar-induced gravitational waves (SIGWs) in different cosmological scenarios. We review three cases for the post-inflationary evolution: the standard case of radiation domination (RD) $c_{s}^{2}=w=1/3$ (see the review [12]), a sudden transition from an early matter-dominated (eMD) phase to RD (see e.g. [19]), and an adiabatic perfect fluid domination with general (constant) equation of state $c_{s}^{2}=w$ [12]. These three cases will then be analysed respectively in sections 4.1, 4.2, and 4.3.
@@
On large cosmological scales, such as those probed by the cosmic microwave background (CMB) with characteristic wavenumbers $k\sim 0.001\mathrm{Mpc}^{-1}$, observations constrain the amplitude of scalar perturbations to be of order $\mathcal{P}_{\zeta}\sim 10^{-9}$ [1]. This leads to an unobservable small SIGW signal on CMB scales. At much smaller scales ($k\gg 1\mathrm{Mpc}^{-1}$), however, there are no direct observational bounds on $\mathcal{P}_{\zeta}$, allowing for the possibility of significant enhancement. In such scenarios, the induced gravitational wave spectrum can attain amplitudes within the sensitivity range of a variety of GW detectors. This opens a new observational window into the small-scale structure of the primordial Universe, complementary to CMB and large-scale structure surveys, which are sensitive to larger scales.
@@
In the standard thermal history, inflation is followed by reheating, which initiates a radiation-dominated (RD) era. Scalar perturbations that exit the horizon during inflation re-enter during RD and source GWs at second order. However, SIGWs are not merely sensitive to the inflationary perturbations, they also offer a probe of the Universe’s post-inflationary thermal history, particularly the epoch between inflation and Big Bang Nucleosynthesis (BBN), which remains largely unconstrained by conventional observations. This period could include non-standard cosmological phases, such as early matter domination or kination, motivated by scenarios in high-energy physics and string theory [^2] See [20] for a recent review on string cosmology, and [21, 22] for reviews on non-standard post-inflationary histories..
The energy density spectrum of SIGWs takes the general form [12]
@@
    Radiation domination: This is the most well-studied case in the literature, see e.g. [12], and corresponds to an equation of state with $w=1/3$. The expression for the induced $\Omega_{\rm GW}$ results
[…]
@@
[…]
@@
    In our work we compute the GW density $\Omega_{\mathrm{GW}}$ for this case using the numerical code SIGWAY[23].
@@
    Early matter domination: We consider a scenario in which a period of early matter domination (eMD) precedes the onset of radiation domination (RD), with a sudden transition between the two phases—referred to here as reheating. A distinctive feature of this setup is the sharp enhancement of the scalar-induced gravitational wave (SIGW) signal immediately after the transition [19]. This enhancement arises because, during eMD, the scalar source term responsible for GW production remains constant and does not efficiently generate GWs. However, the onset of radiation domination leads to efficient GW production. This characteristic imprint in the GW spectrum may fall within the sensitivity reach of upcoming observatories such as LISA, providing a unique observational window into the pre-BBN universe and the reheating dynamics.
@@
    It is important to note that this enhancement is sensitive to the details of the transition. The assumption of a sudden transition corresponds to a change in the background evolution occurring on a timescale much shorter than the Hubble time at reheating. In realistic scenarios, however, the transition is expected to be gradual. During such an intermediate period, the evolution of perturbations is continuous, and their behaviour cannot be fully captured by either the pure eMD or RD approximations. Several studies have explored this regime using smoothed background interpolations [24, 25, 26, 27], showing that the resulting SIGW spectrum is modified in both amplitude and shape compared to the sudden-transition limit.
@@
    As discussed in [19], assuming a sudden transition, however, remains a reasonable approximation in two key regimes: for modes with $k\ll k_{\mathrm{rh}}$, which remain superhorizon at reheating and thus have constant curvature perturbations unaffected by the nature of the transition; and for modes with $k\gg k_{\mathrm{rh}}$, which are already subhorizon and oscillating well before reheating, making their evolution similarly insensitive to how reheating proceeds. The approximation begins to fail, however, for modes with $k\sim k_{\mathrm{rh}}$, which enter the horizon around the time of reheating and are therefore more sensitive to whether the transition is sharp or gradual. That said, if the peak of the primordial power spectrum corresponds to modes that enter the horizon well before reheating, then the dominant contribution to the induced gravitational wave signal comes from earlier times. In this case, reheating does not introduce a significant new source of GWs, and only affects a narrow range of modes near $k_{\mathrm{rh}}$, which do not dominate the overall spectrum. For instance, in models with a sharp infrared (IR) cutoff at some scale $k_{\text{IR }}$, where $k_{\text{IR }}\gg k_{\text{rh }}$, all relevant modes contributing to GW production enter the horizon before reheating. As long as the evolution of induced GWs is also correctly connected across the transition, then we can adopt this sudden transition scenario. This approach is demonstrated by numerical results such as [19] and [23].
@@
    where as before, a subscript ‘rh’ denotes the time of reheating and $x\equiv k\eta$. In sudden reheating scenarios, we focus on gravitational waves induced during the radiation-dominated era by scalar perturbations that previously experienced an early matter-dominated phase on subhorizon scales. While GWs can also be generated during the eMD era [28, 29], their amplitude is typically much smaller than those produced during the RD era, where conditions are more favourable for GW amplification. This is particularly the case in sudden reheating scenarios [19]. Thus, we consider primarily the RD contribution, and modes with $k>k_{\mathrm{rh}}$, which experience a non-trivial evolution before the universe transitioned to radiation domination. The kernel is approximated in terms of two dominant contributions given in terms of the more convenient variables, $s,t$ defined as
@@
    This approximation is valid for large-scale modes with $k\ll k_{\mathrm{cut}}$, where $k_{\mathrm{cut}}$ is the scale corresponding to the onset of eMD, or the scale where density perturbations become non-linear [19, 24].
@@
    In our work, by means of these analytical forms of the kernel, we compute the GW density $\Omega_{\mathrm{GW}}$ using the numerical code SIGWAY[23].
@@
    The gravitational wave (GW) spectrum under such conditions has been studied in [16, 17, 18] (see the review [12] for a detailed discussion). We are interested in particular to the regime where all relevant modes re-enter the horizon well before reheating, i.e., $k\gg k_{\mathrm{rh}}$. The kernel averaged over oscillations is given by:
[…]
@@
[…]
@@
    ×$\times\left\{\left(\mathrm{P}_{b}^{-b}(y)+\frac{b+2}{b+1}\mathrm{P}_{b+2}^{-b}(y)\right)^{2}\Theta\left[c_{s}(u+v)-1\right]\right.$
@@
    In our analysis, we use a modified version of the public code SIGWFast [30] to compute the spectrum $\Omega_{\mathrm{GW}}(f)$ [^3] Where we convert the wavenumber to frequency using ${k}/{\text{Mpc}^{-1}}\simeq 6.5\times 10^{14}f/\mathrm{Hz}.$, given an input curvature power spectrum, a constant equation of state $w$, and a transition (or reheating) scale $f_{\text{rh }}$. The SIGWFast code follows the same conventions which we adopt throughout this work  [^4] Up to a minor redefinition of the variables $u,v$ in terms of $s,d$..
@@ 3.1 Spline Parametrization of the Power Spectrum
The placement of the nodes is not predefined, but it is treated as free parameter to be inferred from the data, ensuring that the model remains adaptable. Only the first and last node are held fixed, based on the range over which the curvature power spectrum is interpolated. This range is typically chosen to be slightly wider than the range over which the $\Omega_{\rm GW}$ is observed. [^5] In our analysis, we find taking $k_{\rm min}=10k_{\rm min,\mathrm{GW}}$ and $k_{\rm max}=10k_{\rm max,\mathrm{GW}}$ to work well for most examples. The free nodes are ordered such that $k_{i-1}<k_{i}<k_{i+1}$, using a bijective ordering transformation that maps a set of variables in $[0,1]$ to an ordered set of variables in the range $[k_{\rm min},k_{\rm max}]$ (see e.g. [31, 32]). Such transformation is characterized by a constant Jacobian, thus if we initially place flat priors on the node locations – as we do here – the ordering transform preserves this flat shape of the prior. For the power spectrum amplitudes at the node positions, we take uniform priors $\log_{10}A_{i}\in[-8,-1]$, unless otherwise stated.
This free-form reconstruction approach allows the data to determine the shape of the power spectrum by adjusting both the positions and amplitudes of the spline nodes. However, increasing model flexibility risks overfitting, which can be mitigated by selecting the optimal number of nodes using Bayesian evidence – naturally balancing model complexity against goodness of fit. Building on this idea, we develop a fully Bayesian framework tailored to this problem. Similar strategies have been employed to reconstruct the primordial curvature perturbations from CMB anisotropies (see e.g., [33, 32] and references therein).

### v3

@@ 1 Introduction
The power spectrum of curvature perturbations, $\mathcal{P}_{\zeta}(k)$, plays a crucial role in understanding the early Universe and the formation of cosmic structure. While the cosmic microwave background (CMB) and large-scale structure of the universe provide constraints on its features at large cosmological scales (see e.g. [Planck:2018vyg]), smaller scales remain relatively unconstrained. One promising avenue to probe this regime is through scalar-induced gravitational waves (SIGWs), which arise as a second-order effect of primordial density perturbations: see e.g. [Matarrese:1992rp, Ananda:2006af, Baumann:2007zm, Saito:2009jt, Bugaev:2009zh, Assadullahi:2009jc, Alabidi:2012ex, Alabidi:2013lya, Espinosa:2018eve, Kohri:2018awv] for original and key papers, and [Domenech:2021ztg] for a comprehensive, detailed review. Such induced gravitational waves carry valuable information about the small-scale features of $\mathcal{P}_{\zeta}(k)$ and may be detected by future gravitational wave observatories, such as the Laser Interferometer Space Antenna (LISA) [LISA:2024hlh] and the Einstein Telescope (ET) [Abac:2025saz]. Enhanced scalar fluctuations, leading to observable GW signals, can be produced on models leading to primordial black holes (PBH), see e.g. [Ozsoy:2023ryl] for a review. In addition to encoding information about the primordial power spectrum, the gravitational wave (GW) spectrum is also sensitive to the expansion history of the Universe (see e.g. [Domenech:2019quo, Hajkarim:2019nbx, Domenech:2020kqm]). This is good news since deviations from a purely radiation-dominated phase at very early times can leave distinct imprints on the shape and amplitude of the GW spectrum, which can be inferred from data. Such departures might arise from exotic physics or alternative cosmological scenarios, and they can be parametrized in terms of an effective equation of state parameter [Domenech:2021ztg]. A detection of SIGWs, therefore, offers a unique opportunity not only to reconstruct $\mathcal{P}_{\zeta}(k)$, but also to infer the background expansion dynamics.
@@ 2 Induced GWs and the Universe’s expansion history
In this section we collect and outline the key equations and conventions relevant for the computation of scalar-induced gravitational waves (SIGWs) in different cosmological scenarios. We review three cases for the post-inflationary evolution: the standard case of radiation domination (RD) $c_{s}^{2}=w=1/3$ (see the review [Domenech:2021ztg]), a sudden transition from an early matter-dominated (eMD) phase to RD (see e.g. [inomata_enhancement_2019]), and an adiabatic perfect fluid domination with general (constant) equation of state $c_{s}^{2}=w$ [Domenech:2021ztg]. These three cases will then be analysed respectively in sections 4.1, 4.2, and 4.3.
@@
On large cosmological scales, such as those probed by the cosmic microwave background (CMB) with characteristic wavenumbers $k\sim 0.001\mathrm{Mpc}^{-1}$, observations constrain the amplitude of scalar perturbations to be of order $\mathcal{P}_{\zeta}\sim 10^{-9}$ [Planck:2018vyg]. This leads to an unobservable small SIGW signal on CMB scales. At much smaller scales ($k\gg 1\mathrm{Mpc}^{-1}$), however, there are no direct observational bounds on $\mathcal{P}_{\zeta}$, allowing for the possibility of significant enhancement. In such scenarios, the induced gravitational wave spectrum can attain amplitudes within the sensitivity range of a variety of GW detectors. This opens a new observational window into the small-scale structure of the primordial Universe, complementary to CMB and large-scale structure surveys, which are sensitive to larger scales.
@@
In the standard thermal history, inflation is followed by reheating, which initiates a radiation-dominated (RD) era. Scalar perturbations that exit the horizon during inflation re-enter during RD and source GWs at second order. However, SIGWs are not merely sensitive to the inflationary perturbations, they also offer a probe of the Universe’s post-inflationary thermal history, particularly the epoch between inflation and Big Bang Nucleosynthesis (BBN), which remains largely unconstrained by conventional observations. This period could include non-standard cosmological phases, such as early matter domination or kination, motivated by scenarios in high-energy physics and string theory [^2] See [Cicoli:2023opf] for a recent review on string cosmology, and [Allahverdi:2020bys, Batell:2024dsi] for reviews on non-standard post-inflationary histories..
The energy density spectrum of SIGWs takes the general form [Domenech:2021ztg]
@@
    Radiation domination: This is the most well-studied case in the literature, see e.g. [Domenech:2021ztg], and corresponds to an equation of state with $w=1/3$. The expression for the induced $\Omega_{\rm GW}$ results
    $\Omega_{\mathrm{GW},\mathrm{RD}}=\int_{0}^{\infty}dv\int_{|1-v|}^{1+v}du\mathcal{T}_{\rm RD}\left(u,v,c_{s}\right)\mathcal{P}_{\zeta}(ku)\mathcal{P}_{\zeta}(kv)\,$, (2.5)
@@
    $\times\left(\frac{\pi}{2}\,4\,y^{2}\,\Theta\!\bigl[c_{s}(u+v)-1\bigr]+\Bigl(1-\tfrac{1}{2}\,y\ln\frac{1+y}{1-y}\Bigr)^{2}\right)\,$,
@@
    In our work we compute the GW density $\Omega_{\mathrm{GW}}$ for this case using the numerical code SIGWAY[LISACosmologyWorkingGroup:2025vdz].
@@
    Early matter domination: We consider a scenario in which a period of early matter domination (eMD) precedes the onset of radiation domination (RD), with a sudden transition between the two phases—referred to here as reheating. A distinctive feature of this setup is the sharp enhancement of the scalar-induced gravitational wave (SIGW) signal immediately after the transition [inomata_enhancement_2019]. This enhancement arises because, during eMD, the scalar source term responsible for GW production remains constant and does not efficiently generate GWs. However, the onset of radiation domination leads to efficient GW production. This characteristic imprint in the GW spectrum may fall within the sensitivity reach of upcoming observatories such as LISA, providing a unique observational window into the pre-BBN universe and the reheating dynamics.
@@
    It is important to note that this enhancement is sensitive to the details of the transition. The assumption of a sudden transition corresponds to a change in the background evolution occurring on a timescale much shorter than the Hubble time at reheating. In realistic scenarios, however, the transition is expected to be gradual. During such an intermediate period, the evolution of perturbations is continuous, and their behaviour cannot be fully captured by either the pure eMD or RD approximations. Several studies have explored this regime using smoothed background interpolations [Inomata_gradual_2019, pearce2024InterpolatingFastandSlow, pearce2025EarlyMatterDominatedEpochs, Kumar:2024hsi], showing that the resulting SIGW spectrum is modified in both amplitude and shape compared to the sudden-transition limit.
@@
    As discussed in [inomata_enhancement_2019], assuming a sudden transition, however, remains a reasonable approximation in two key regimes: for modes with $k\ll k_{\mathrm{rh}}$, which remain superhorizon at reheating and thus have constant curvature perturbations unaffected by the nature of the transition; and for modes with $k\gg k_{\mathrm{rh}}$, which are already subhorizon and oscillating well before reheating, making their evolution similarly insensitive to how reheating proceeds. The approximation begins to fail, however, for modes with $k\sim k_{\mathrm{rh}}$, which enter the horizon around the time of reheating and are therefore more sensitive to whether the transition is sharp or gradual. That said, if the peak of the primordial power spectrum corresponds to modes that enter the horizon well before reheating, then the dominant contribution to the induced gravitational wave signal comes from earlier times. In this case, reheating does not introduce a significant new source of GWs, and only affects a narrow range of modes near $k_{\mathrm{rh}}$, which do not dominate the overall spectrum. For instance, in models with a sharp infrared (IR) cutoff at some scale $k_{\text{IR }}$, where $k_{\text{IR }}\gg k_{\text{rh }}$, all relevant modes contributing to GW production enter the horizon before reheating. As long as the evolution of induced GWs is also correctly connected across the transition, then we can adopt this sudden transition scenario. This approach is demonstrated by numerical results such as [inomata_enhancement_2019] and [LISACosmologyWorkingGroup:2025vdz].
@@
    where as before, a subscript ‘rh’ denotes the time of reheating and $x\equiv k\eta$. In sudden reheating scenarios, we focus on gravitational waves induced during the radiation-dominated era by scalar perturbations that previously experienced an early matter-dominated phase on subhorizon scales. While GWs can also be generated during the eMD era [Alabidi_2013, Assadullahi_2009], their amplitude is typically much smaller than those produced during the RD era, where conditions are more favourable for GW amplification. This is particularly the case in sudden reheating scenarios [inomata_enhancement_2019]. Thus, we consider primarily the RD contribution, and modes with $k>k_{\mathrm{rh}}$, which experience a non-trivial evolution before the universe transitioned to radiation domination. The kernel is approximated in terms of two dominant contributions given in terms of the more convenient variables, $s,t$ defined as
@@
    This approximation is valid for large-scale modes with $k\ll k_{\mathrm{cut}}$, where $k_{\mathrm{cut}}$ is the scale corresponding to the onset of eMD, or the scale where density perturbations become non-linear [inomata_enhancement_2019, Inomata_gradual_2019].
@@
    In our work, by means of these analytical forms of the kernel, we compute the GW density $\Omega_{\mathrm{GW}}$ using the numerical code SIGWAY[LISACosmologyWorkingGroup:2025vdz].
@@
    The gravitational wave (GW) spectrum under such conditions has been studied in [Domenech:2019quo, Hajkarim:2019nbx, Domenech:2020kqm] (see the review [Domenech:2021ztg] for a detailed discussion). We are interested in particular to the regime where all relevant modes re-enter the horizon well before reheating, i.e., $k\gg k_{\mathrm{rh}}$. The kernel averaged over oscillations is given by:
    $\begin{aligned} \overline{I^{2}(u,v,x)}=&x^{-2(b+1)}4^{2b}\Gamma^{4}[b+3/2]\left(\frac{2b+3}{b+2}\right)^{2}\frac{\left|1-y^{2}\right|^{b}}{2c_{s}^{4}u^{2}v^{2}}\\ &\times\left\{\left(\mathrm{P}_{b}^{-b}(y)+\frac{b+2}{b+1}\mathrm{P}_{b+2}^{-b}(y)\right)^{2}\Theta\left[c_{s}(u+v)-1\right]\right.\\ &+\frac{4}{\pi^{2}}\left(\mathrm{Q}_{b}^{-b}(y)+\frac{b+2}{b+1}\mathrm{Q}_{b+2}^{-b}(y)\right)^{2}\Theta\left[c_{s}(u+v)-1\right]\\ &\left.+\frac{4}{\pi^{2}}\left(\mathcal{Q}_{b}^{-b}(-y)+2\frac{b+2}{b+1}\mathcal{Q}_{b+2}^{-b}(-y)\right)^{2}\Theta\left[1-c_{s}(u+v)\right]\right\}\,.\end{aligned}$ (2.13)
@@
    $\Omega_{\mathrm{GW},\mathrm{rh}}=\left(\frac{k}{k_{\mathrm{rh}}}\right)^{-2b}\int_{0}^{\infty}dv\int_{|1-v|}^{1+v}du\mathcal{T}\left(u,v,b,c_{s}\right)\mathcal{P}_{\zeta}(ku)\mathcal{P}_{\zeta}(kv)$ (2.17)
@@
    $\times\left\{\left(\mathrm{P}_{b}^{-b}(y)+\frac{b+2}{b+1}\mathrm{P}_{b+2}^{-b}(y)\right)^{2}\Theta\left[c_{s}(u+v)-1\right]\right$.
@@
    In our analysis, we use a modified version of the public code SIGWFast [Witkowski:2022mtg] to compute the spectrum $\Omega_{\mathrm{GW}}(f)$ [^3] Where we convert the wavenumber to frequency using ${k}/{\text{Mpc}^{-1}}\simeq 6.5\times 10^{14}f/\mathrm{Hz}.$, given an input curvature power spectrum, a constant equation of state $w$, and a transition (or reheating) scale $f_{\text{rh }}$. The SIGWFast code follows the same conventions which we adopt throughout this work  [^4] Up to a minor redefinition of the variables $u,v$ in terms of $s,d$..
@@ 3.1 Spline Parametrization of the Power Spectrum
The placement of the nodes is not predefined, but it is treated as free parameter to be inferred from the data, ensuring that the model remains adaptable. Only the first and last node are held fixed, based on the range over which the curvature power spectrum is interpolated. This range is typically chosen to be slightly wider than the range over which the $\Omega_{\rm GW}$ is observed. [^5] In our analysis, we find taking $k_{\rm min}=10k_{\rm min,\mathrm{GW}}$ and $k_{\rm max}=10k_{\rm max,\mathrm{GW}}$ to work well for most examples. The free nodes are ordered such that $k_{i-1}<k_{i}<k_{i+1}$, using a bijective ordering transformation that maps a set of variables in $[0,1]$ to an ordered set of variables in the range $k_{\rm min},k_{\rm max}]$ (see e.g. [Buscicchio_prior:2019, Handley:2019fll]). Such transformation is characterized by a constant Jacobian, thus if we initially place flat priors on the node locations – as we do here – the ordering transform preserves this flat shape of the prior. For the power spectrum amplitudes at the node positions, we take uniform priors $\log_{10}A_{i}\in[-8,-1]$, unless otherwise stated.
This free-form reconstruction approach allows the data to determine the shape of the power spectrum by adjusting both the positions and amplitudes of the spline nodes. However, increasing model flexibility risks overfitting, which can be mitigated by selecting the optimal number of nodes using Bayesian evidence – naturally balancing model complexity against goodness of fit. Building on this idea, we develop a fully Bayesian framework tailored to this problem. Similar strategies have been employed to reconstruct the primordial curvature perturbations from CMB anisotropies (see e.g., [Planck:2018jri, Handley:2019fll] and references therein).